Computer Geeks has sales of $566980, a profit margin of 0.53, a total asset turnover rate of 1.13, and an equity multiplier of 1.39. What is the return on equity?
Answer #1

Net Income = Sales * Profit margin

= 566980*0.53

= 300499.40

Total assets = Sales/TAT = 566980/1.13 = 501752.2

Hence

Equity Multiplier = Total assets/ Equity

Equity = 501752.2/1.39 = 360972.8

ROE = Net Income/ Equity

= 300499.4/ 360972.8

=83.25%